OpenStudy (anonymous):

Russel throws a soda from the balcony of his house to his friend Jones in the pool. If the soda is released 20 feet from the ground, and is thrown with a downward velocity of 2 ft/s, how long till the soda hits the pool, ignoring air resitance? Use the formula 2003-19-06-00-00_files/i0180000.jpg, where h is the height of the soda in feet and t is the time in seconds since Russel threw it. Round your answer to the nearest tenth.

OpenStudy (welshfella):

use one of the equations of motion for constant acceleration height s = 20 ft , initial velocity u = 2 ft s-1, acceleration due to gravity= 32 ft s-2 and time t = ? s = ut + 0.5at^2 20 = 2t + 16t^2 16t^2 + 2t - 20 = 0 solve for t